OpenStudy (anonymous):

Hey can someone help me with a couple of questions? Simplify 1. sqrt 4/ 2a 2. ^3 sqrt 6b/ 9a^2 3. sqrt5/ sqrt3 - sqrt2 4. sqrt5 x sqrt4 Show step by step

OpenStudy (anonymous):

\[\frac{ \sqrt{4} }{ 2a }\] \[\frac{ 2 }{ 2a }\]

OpenStudy (solomonzelman):

Hami, you want us to help you, not to do all your work, correct ?

OpenStudy (anonymous):

I actually got the answers, thanks though!